OpenStudy (anonymous):

divide and simplify x^2-25/x divided by x-5/x+2= type a fraction and leave your answer in factored form

OpenStudy (anonymous):

When ever you divide by a fraction you can instead multiply by the reciprocal. As what I understand from your question you wrote: \[[ (x^2 -25)/x ]/ [(x-5)/(x+2)]\] This changes to: \[[ (x^2 -25)/x ]*[(x+2)/(x-5)]\] Now there is something cool we can do... like factor that nasty x^2-25 We should get this when we factor: \[[ (x-5)(x+5)/x ]*[(x+2)/(x-5)]\] Which when we multiply across we can see the x-5 cancel out: \[[(x+5)(x+2)]/x\] And we are done!

OpenStudy (anonymous):

oreostar, try the \frac{numerator}{denominator} macro. Like:\[\frac{\frac{x^2-25}{x}}{\frac{x-5}{x+2}}\]which is written with \frac{\frac{x^2-25}{x}}{\frac{x-5}{x+2}}

OpenStudy (anonymous):

Cool, I looked in the equation thing for it and couldn't find a easy button. I will use this in the future. It's so ugly written out the way I did it.